9s=6a
s=2x
a=(x+.24)
Find the value of each variable.

Answer by stanbon(48545)

9s=6a
s=2x
a=(x+.24)
Find the value of each variable.
-----------
Substitute for s and for a to get:
9(2x) = 6(x+0.24)
18x = 6x + 1.44
12x = 1.44
x = 0.12
Therefore
s=2x = 0.24
a=(x+0.24) = 0.12+0.24 = 0.36
================
Cheers,
Stan H.